OpenStudy (anonymous):

what are the factors of x^2-81?

OpenStudy (amistre64):

add 81 to each sides and sqrt it that gives you the roots and the factors can be reconstructed from the roots if need be

OpenStudy (amistre64):

x^2=81 x = +- 9 (x-9)(x+9)

OpenStudy (anonymous):

= (x)^2 - (+-9)^2 =(x+9)(x-9) Ans